# Sparse Hanson-Wright Inequalities with Applications

### v2

@@ 1 Introduction
Let $X=(X_{1},\cdots,X_{n})^{T}$ be a random vector with independent centered sparse $\alpha$-subexponential entries, i.e., $X_{i}=x_{i}\xi_{i}$ where $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p_{i})$ and $\xi_{i}$ is a centered random variables with $\|\xi_{i}\|_{L_{r}}\leq Kr^{\frac{1}{\alpha}}$ for any $r\geq 1$ and some $\alpha>0$. Let $A$ be an $n\times n$ matrix. Then for any $t>0$,
@@
$\leq$ $C^{\prime}\exp\left(-C_{\alpha}\min\left\{\frac{t^{2}}{K^{4}(\sum_{i}a_{ii}^{2}p_{i}+\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j})},\frac{t}{K^{2}\gamma_{1,\infty}},\left(\frac{t}{K^{2}\|A\|_{\max}}\right)^{\min\{\frac{\alpha}{2},\frac{1}{2}\}}\right\}\right),$
where
Note that the definition of $\gamma_{1,\infty}$ in (1.8) is consistent with the diagonal-free case in (1.3). Furthermore, Theorem 2, Corollary 1, and Corollary 2 can be extended through generalizations similar to that of Theorem 1, as discussed in Remark 1.1. We omit the details for brevity.
@@
The proof of Hanson-Wright inequalities for sub-gaussian random vectors in [35] replies on bounding the exponential generating function with a Gaussian comparison argument. Due to the sparsity parameters, a direct adaptation of the proof in [35] will yield a dependence on the sub-gaussian norm of $x_{i}\xi_{i}$, which does not have the optimal dependence on $p_{i}$. Moreover, this technique does not apply to $\alpha$-subexponential random variables. Instead, our proof is based on the moment method, which captures the correct variance term in the sub-gaussian tail part and the sub-exponential tail part also depends on the sparsity parameters. The moment method approach was used in [24, 37, 38]. Our inequality improves the one in [24] to have sub-gaussian and sub-exponential tails, and generalize the inequalities from [37] to sparse $\alpha$-subexponential random vectors with a simpler proof.
We present two applications of the new sparse Hanson-Wright inequality. The first application is a new local law and eigenvector delocalization for Hermitian random matrices with independent sparse $\alpha$-subexponential entries. The second is the concentration of the norm of the linear transformation of an $\alpha$-sparse subexponential random vector.
@@
Organization: The rest of the paper is organized as follows. Section 2 reviews existing literature on sparse Hanson-Wright inequalities, along with a discussion and comparison of our findings with previous work. We present two applications of our sparse Hanson-Wright inequality in Section 3. The proofs of our main theoretical results are provided in Sections 4 and 5, while Section 6 contains the proofs of the applications. Additional technical results and generalizations of our theorems are presented in Appendices A and B, respectively. Finally, Appendices C and D provide the proofs of auxiliary results used in the applications.
@@ 2 Previous results and comparison
Recent progress in the Hanson-Wright-type inequalities and more generally, the polynomials of independent random variables, can be found in [42, 2, 47, 24, 46, 27, 48, 31, 15, 22, 36, 6, 26, 38, 16, 36, 22, 13, 43] and references therein. Specifically, an extension to $\alpha$-subexponential random vectors was obtained in [22]:
THEOREM 4 (Corollary 1.4 (2) in [22]).
For any $\alpha\in(0,1]$, let $X_{1},\dots,X_{n}$ be independent centered random variables with $\|X_{i}\|_{\Psi_{\alpha}}\leq K$, and $A$ be a symmetric matrix. Then for any $t\geq 0$,
@@
The extension of (2.2) to any $\alpha\in(0,2]$ is discussed in Sambale’s work [36].
The concentration of multivariate polynomials of independent random variables has been a subject of study since the work of Kim and Vu [26]. Schudy and Sviridenko [37] derived concentration inequalities for polynomials of independent central moment bounded random variables, improving upon their earlier findings in [38]. A random variable $Z$ is considered central moment bounded if $\mathbb{E}(|Z-\mathbb{E}Z|^{i})\leq iL\mathbb{E}(|Z-\mathbb{E}Z|^{i-1})$ for any integer $i\geq 1$. This class of central moment bounded random variables encompasses bounded and log-concave random variables, with Gaussian and exponential distributions as special examples.
The results by Schudy and Sviridenko [37], when applied specifically to quadratic forms of sparse exponential random variables, yield the following:
THEOREM 5 (Special case of Theorem 1.3 in [37]).
Let $X_{1},\cdots,X_{n}$ be independent centered sparse exponential random variables. Each $X_{i}=x_{i}\eta_{i}$ where $x_{i},\eta_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p_{i})$ and $\eta_{i}$ is a centered $\mathrm{Exp}(K^{-1})$ random variable. Let $A$ be an $n\times n$ symmetric, diagonal-free matrix. For any $t>0$,
@@
This result aligns with our Theorem 1 for the sparse exponential random variables. However, our findings extend beyond this, encompassing sparse sub-exponential random variables and further generalizing to sparse $\alpha$-subexponential random variables, as given in Theorem 2. Our proof is also simpler compared to [38]. It is worth noting that, similar to the discussion in [22, Section 1.2], a central moment bounded random variable $Z$ can be shown to be sub-exponential, i.e., $\|Z\|_{\Psi_{1}}<\infty$. The converse implication is not immediately evident and may not hold in all cases.
Another line of research focuses on the tail bounds of (unbounded) U-statistics. Chakrabortty and Kuchibhotla [14] examine a particular class of U-statistics. These U-statistics take the form: $U_{n}=\sum_{1\leq i\neq j\leq n}\phi(Y_{i})w_{ij}(X_{i},X_{j})\phi(Y_{j})$, where $Z_{1}=(X_{1},Y_{1}),\cdots,Z_{n}=(X_{n},Y_{n})$ are independent random vectors. When applied to sparse $\alpha$-subexponential random variables, their results yield the following moment estimates:
THEOREM 6 (Theorem 1 from [14]).
Under the same assumptions as Theorem 2, denote $\alpha^{*}:=\min\{\alpha,1\}$. For any integer $r\geq 1$,
$\|X^{T}AX\|_{L^{r}}\lesssim$ $\sqrt{r}\cdot K_{\alpha}^{2}p\|A\|_{F}+r\cdot K_{\alpha}^{2}p\|A\|$
$+r^{\frac{1}{\alpha^{*}}}\cdot K_{\alpha}^{2}\left[(\log n)^{\frac{1}{2}+\frac{1}{\alpha}}\sqrt{p}\|A\|_{2,\infty}+(\log n)^{1+\frac{2}{\alpha}}\|A\|_{\max}\right]$
$+r^{\frac{1}{2}+\frac{1}{\alpha^{*}}}\cdot K_{\alpha}^{2}(\log n)^{\frac{1}{\alpha}}\sqrt{p}\|A\|_{2,\infty}+r^{\frac{2}{\alpha^{*}}}\cdot K_{\alpha}^{2}(\log n)^{\frac{2}{\alpha}}\|A\|_{\max}.$ (2.4)
The bound above comprises a mixture of four terms. Notably, for $\alpha\in(0,1]$, the tail bound implied by these moment estimates is consistent with the result in Theorem 4 for constant $p$, up to the $\log n$ factors.
An intriguing future direction would be to derive a refined concentration inequality that encompasses both our result (Theorem 2) and the above results. Recent work by [7] has extended the results in [14], generalizing their findings to a broader class of unbounded, heavy-tailed U-statistics.
There is also considerable interest in the study of quadratic forms involving sparse sub-gaussian random variables [47, 6, 34]. A key result was obtained by Zhou [47]:
THEOREM 7 (Theorem 1.1 from [47]).
Let $X_{1},\cdots,X_{n}$ be independent centered sparse sub-gaussian random variables, i.e. $X_{i}=x_{i}\xi_{i}$ where $x_{i},\xi_{i}$ are independent, $x_{i}\sim\mathrm{Ber}(p_{i})$ and $\xi_{i}$ is a centered sub-gaussian random variables with $\|\xi_{i}\|_{\Psi_{2}}\leq K$. Let $A$ be an $n\times n$ matrix. Denote $S_{n}=\sum_{i,j}a_{ij}x_{i}x_{j}\xi_{i}\xi_{j}$. Then for any $t>0$,
$\mathbb{P}\left(\left|S_{n}-\mathbb{E}(S_{n})\right|>t\right)\leq 2\exp\left(-c\min\left\{\frac{t^{2}}{K^{4}(\sum_{k=1}^{n}p_{k}a_{kk}^{2}+\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j})},\frac{t}{K^{2}\|A\|}\right\}\right).$
Our Corollary 2, when applied to $\alpha=2$, reveals up to three different regimes, in contrast to the two regimes as above. A comparable bound to [47] is derived in [34] for the generalized setting $S=\sum_{i,j}a_{ij}x_{1i}x_{2j}\xi_{1i}\xi_{2j}$ where the pairs $(x_{1i},x_{2i})$ and $(\xi_{1i},\xi_{2i})$ are independent, although $x_{1i},x_{2i}$ or $\xi_{1i},\xi_{2i}$ may be correlated. Furthermore, [6] has recently established a result similar to [47], under the assumptions that $\|A\|\leq 1$ and all Bernoulli random variables share the same parameter $p$. The concentration inequality in Lemma 2.8 from [6] yields a sharp bound assuming $1\gg p\gg\log n/n$.
The concentration of quadratic forms involving (sparse) Bernoulli random variables, specifically $\sum_{i,j}a_{ij}x_{i}x_{j}$ where $x_{i}$’s are independent Bernoulli (or centered Bernoulli) random variables, has been a subject of extensive study in the literature. While not the primary focus of our current work, we find it valuable to offer a concise summary of selected results in this area.
For simplicity, for the rest of this section, we assume that $A$ is symmetric, diagonal-free, and $x_{i}$ are independent centered Bernoulli random variables with parameter $p_{i}$. Hence, $\sum_{i\neq j}a_{ij}x_{i}x_{j}:=x^{T}Ax$ is centered.
Latała [29] provides precise moment estimates for the quadratic forms of Rademacher random variables, where all $p_{i}=\frac{1}{2}$. For any $r\geq 1$,
$C^{-1}(\|A\|_{\{1,2\},r}+\|A\|_{\{1\},\{2\},r})\leq\|x^{T}Ax\|_{L^{r}}\leq C(\|A\|_{\{1,2\},r}+\|A\|_{\{1\},\{2\},r}).$ (2.5)
The definitions of matrix norms above are involved and can be found in [29] or [3, Section 5.2]. These norms satisfy that $\|A\|_{\{1,2\},r}\leq\sqrt{r}\|A\|_{F}$ and $\|A\|_{\{1\},\{2\},r}\leq r\|A\|$, which allows this result to recover the Hanson-Wright inequality (1.1).
Giné, Latała, and Zine’s study [21] on the concentration of bounded U-statistics offers insights that are useful for deriving tail estimates of random quadratic forms. By applying decoupling results and performing straightforward calculations, one can obtain the following tail estimate as a special case of their Theorem 3.3 in [21]: for any $t>0$,
where $B=(\max_{i}p_{i})\|A\|$ and $D=\max_{j}\sqrt{\sum_{i}a_{ij}^{2}p_{i}}$. The $t^{2}$ and $t^{1/2}$ terms in (2.6) match our Theorem 1 for bounded random variables. The $t$ term in some cases (e.g., all $p_{i}$ are the same) can be better than Theorem 1, while Theorem 1 does not have the $t^{2/3}$ term.
The result by Schudy and Sviridenko [37], as mentioned earlier, applies to the sparse Bernoulli random variable case. Their findings yield the tail bound:
$\mathbb{P}(|x^{T}Ax|\geq t)\leq 2\exp\left(-C\min\left\{\frac{t^{2}}{\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j}},\frac{t}{\max_{i}\sum_{j}|a_{ij}|p_{j}},\left(\frac{t}{\|A\|_{\max}}\right)^{1/2}\right\}\right).$ (2.7)
Note that a direct adaptation of our proof for Theorem 1 provides the same bound under this setting.
Bellec [8] derived a sharper version of the Hanson-Wright inequality than (1.1) for random variables satisfying a Bernstein moment condition. This improved inequality captures the variance of $X^{T}AX$ in the sub-gaussian tail. Their result applies to sparse Bernoulli random variables. He, Knowles, and Marcozzi [24] established an upper bound on the $L_{r}$ norms $\|x^{T}Ax\|_{L^{r}}$, assuming all $p_{i}=p$. Notably, their bound involves $\|A\|_{2,\infty}$ and $\|A\|_{\max}$. This estimate is crucial in analyzing the spectral statistics of sparse Erdős-Rényi random graphs $G(n,p)$ at the critical scale $pn=O(\log n)$. Zhou [48] (see Theorem 2.5) examined the quadratic forms of non-identical independent Bernoulli random variables, providing bounds on the moment generating function (MGF) of the off-diagonal term.

### v3

@@ 1.2 Bennett-type inequalities for sparse random vectors
Let $X=(X_{1},X_{2},\ldots,X_{n})^{T}$ be a random vector with independent centered sparse $\alpha$-subexponential components, where each $X_{i}$ satisfies $\mathbb{E}(|X_{i}|^{r})\leq p_{i}(Kr^{\frac{1}{\alpha}})^{r}$ for all $r\geq 1$ and some $\alpha\in(1,\infty]$.
  • (Linear form) Let $a=(a_{1},\ldots,a_{n})$ be a deterministic vector in $\mathbb{R}^{n}$. Denote $\lambda^{2}={\sum_{i=1}^{n}a_{i}^{2}p_{i}}$. For any $t\gtrsim K\lambda^{2}/\|a\|_{\infty}$,
    $\mathbb{P}\left(\left|\sum_{i=1}^{n}a_{i}X_{i}\right|>t\right)\leq e^{2}\exp\left(-\frac{C_{\alpha}t}{K\|a\|_{\infty}}\left[\log\Big{(}\frac{t\|a\|_{\infty}}{c_{\alpha}K\lambda^{2}}\Big{)}\right]^{1-\frac{1}{\alpha}}\right).$ (1.6)
  • (Quadratic form) Let $A$ be a diagonal-free $n\times n$ matrix. Denote $\sigma^{2}=\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j}$ and $\gamma_{1,\infty}$ as in (1.4). Set $\Lambda:=\max\{\gamma_{1,\infty},(\|A\|_{\max}\sigma^{2})^{\frac{1}{3}}\}$. For any $t\gtrsim K^{2}{\Lambda^{2}}/{\|A\|_{\max}}$,
    $\mathbb{P}\left(\left|X^{T}AX\right|>t\right)\leq e^{2}\exp\left(-C_{\alpha}\sqrt{\frac{t}{K^{2}\|A\|_{\max}}}\left[\log\left(\frac{t\|A\|_{\max}}{c_{\alpha}K^{2}\Lambda^{2}}\right)\right]^{1-\frac{1}{\alpha}}\right).$
Compared to Theorems 1 and 2, when $\alpha>1$, for sufficiently large $t$, the tail probability gains an additional $\log^{1-1/\alpha}(t)$ factor, analogous to the improvement from Bernstein’s inequality to Bennett’s inequality. Notably, for the sparse bounded random variables ($\alpha=\infty$), (1.6) matches Bennett’s inequality in the large deviation regime (up to constants). A detailed comparison is presented in Section 2. The tail behavior described by Bennett’s inequality in the large deviation regime for bounded random variables is known to be sharp (up to universal constants); see [48, Example 2.4]. To the best of our knowledge, these Bennett-type inequalities are new for both the linear form when $\alpha\in(1,\infty)$ and the quadratic form when $\alpha\in(1,\infty]$ of sparse $\alpha$-subexponential random variables.
Before presenting improved precise tail bounds for the linear and quadratic forms for $\alpha$-subexponential random variables with $\alpha\in(1,\infty]$, we introduce several essential notations.
For real numbers $x$ and $y$, when $x\geq-e^{-1}$, the Lambert W function $W(x)$ is defined as the solution to the equation $ye^{y}=x$. When $x\in(-e^{-1},0)$, the equation admits two real solutions, corresponding to two branches of the Lambert W function: the principal branch $W_{0}(x)$ and the lower branch $W_{-1}(x)$. On this interval, $W_{0}(x)$ is an increasing function satisfying $W_{0}(x)\geq-1$, while $W_{-1}(x)$ is a decreasing function satisfying $W_{-1}(x)\leq-1$ (see Figure 1). We compile the basic properties of this special function in Appendix A.2; for a detailed discussion of this function, we refer to [23].
Throughout the context, we denote
$\beta:=1-\frac{1}{\alpha}\quad\text{for }\alpha\in(1,\infty].$ (1.7)
THEOREM 4 (Tail bound for the linear form: $\alpha>1$).
Under the assumptions in Theorem 2 for $\alpha\in(1,\infty]$ where ${\sum_{i=1}^{n}a_{i}^{2}p_{i}}\leq\lambda^{2}$ and $\|a\|_{\infty}\leq M$, we have for any $t>0$,
$\mathbb{P}\left(\left|\sum_{i=1}^{n}a_{i}X_{i}\right|>t\right)\leq e^{2}\exp\left(-C_{\alpha}\min\left\{\frac{t^{2}}{K^{2}\lambda^{2}},\frac{t}{KM}\max\left\{1,H_{\alpha}(T)\right\}\right\}\right),$ (1.8)
where we denote
$H_{\alpha}(T):=\left[-\beta W_{-1}\Big{(}-\frac{1}{\beta}T^{-\frac{1}{\beta}}\Big{)}\right]^{\beta}\mathbf{1}(T>(e/\beta)^{\beta})\quad\text{with}\quad T:=\frac{tM}{c_{\alpha}K\lambda^{2}}.$
REMARK 1.1.
Compared to Theorem 2, the large deviation regime exhibits a new tail bound of the form
$\exp\left(-C_{\alpha}\frac{t}{KM}\cdot H_{\alpha}(T)\right).$
The function $H_{\alpha}(t)$ is increasing in $t$ and grows to infinity as $t\to\infty$. Using the asymptotic expansion of the Lambert W function $W_{-1}(x)\approx\log(-x)-\log(-\log(-x))$ as $x\to 0^{-}$ (see [23]), we obtain that for large $t$,
$H_{\alpha}(T)\approx\left[\log(T)+\beta\log\left(\beta^{-1}\log(T)-\log(\beta^{-1})\right)-\beta\log(\beta^{-1})\right]^{\beta}.$
In the special case where $\alpha=\infty$ and $\beta=1$, this simplifies to
$H_{\infty}(T)\approx\log(T)+\log(\log(T))=\log\left(\frac{tM}{CK\lambda^{2}}\right)+\log\left(\log\left(\frac{tM}{CK\lambda^{2}}\right)\right).$
Next, we present the improved tail bound for the quadratic forms.
THEOREM 5 (Tail bound for the quadratic form: $\alpha>1$).
Under the assumptions in Theorem 1 for $\alpha\in(1,\infty]$ where $\sum_{i\neq j}a_{ij}^{2}p_{i}p_{j}\leq\sigma^{2}$, $\gamma_{1,\infty}\leq\Gamma_{1,\infty}$, and $\|A\|_{\max}\leq M$, we have for any $t>0$,
$\mathbb{P}\left(\left|\sum_{i\neq j}a_{ij}X_{i}X_{j}\right|>t\right)$ (1.9)
@@
$\leq e^{2}\exp\left(-C_{\alpha}\min\left\{\frac{t^{2}}{K^{4}\sigma^{2}},\frac{t}{K^{2}\Gamma_{1,\infty}},\sqrt{\frac{t}{K^{2}M}}\max\left\{1,\min_{i=1,2}H_{\alpha}(T_{i})\right\}\right\}\right),$ (1.10)
where we denote
$H_{\alpha}(T):=\left[-\beta W_{-1}\Big{(}-\frac{1}{\beta}T^{-\frac{1}{\beta}}\Big{)}\right]^{\beta}\mathbf{1}(T>(e/\beta)^{\beta})$
with
$T_{1}:=\frac{M}{\Gamma_{1,\infty}}\left(\frac{t}{c_{\alpha}K^{2}M}\right)^{1/2}\quad\text{and}\quad T_{2}:=\left(\frac{M}{\sigma}\right)^{2/3}\left(\frac{t}{c_{\alpha}^{\prime}K^{2}M}\right)^{1/2}.$
REMARK 1.2.
Due to the implicit nature of the Lambert W function, for practical applications of Theorems 4 and 5, we utilize the bounds established in [21]:
$-1-\sqrt{2u}-u<W_{-1}(-e^{-u-1})<-1-\sqrt{2u}-\frac{2}{3}u$ (1.11)
for $u>0$. One can readily verify that Theorem 3 follows directly from applying these bounds (1.11) to Theorems 4 and 5, observing that $W_{-1}\big{(}-\frac{1}{\beta}T^{-\frac{1}{\beta}}\big{)}=W_{-1}\big{(}-e^{-\log\big{(}\frac{\beta}{e}T^{\frac{1}{\beta}}\big{)}-1}\big{)}$.
@@
The proof of Hanson-Wright inequalities for sub-gaussian random vectors in [52] replies on bounding the exponential generating function with a Gaussian comparison argument. Due to the sparsity parameters, a direct adaptation of the proof in [52] will yield a dependence on the sub-gaussian norm of $X_{i}$, which does not have the optimal dependence on $p_{i}$. Moreover, this technique does not apply to $\alpha$-subexponential random variables. Instead, our proof is based on the moment method, which captures the correct variance term in the sub-gaussian tail part and the subexponential tail part also depends on the sparsity parameters. The moment method approach was used in [34, 54, 55]. Our inequality improves the ones in [34] (corresponding to $\alpha=\infty$) to have Bennett-type inequalities for both linear and quadratic forms given in Theorems 4 and 5. We also generalize and improve the inequalities from [54] to sparse $\alpha$-subexponential random vectors for all the regime $\alpha\in(1,\infty]$. A detailed discussion and comparison with the existing literature is presented in Section 2.
In Section 3, we present two applications of the new sparse Hanson-Wright inequalities. The first application establishes a new local law and eigenvector delocalization for Hermitian random matrices with independent sparse $\alpha$-subexponential entries. Notably, we derive the sharpest known upper bound for the infinity norm of eigenvectors in Wigner matrices with sparse $\alpha$-subexponential entries for any $\alpha\in(0,\infty]$ when the sparsity parameter $p\gtrsim\frac{\log^{2\vee(1+2/\alpha)}(n)}{n}$. For $\alpha\geq 2$, we extend the local law and eigenvector delocalization results to the super-critical regime $p\gtrsim\frac{\log(n)}{n}$, matching the bounds in [34] ($\alpha=\infty$) and generalizing them to any $\alpha\geq 2$ beyond the bounded support case. The second application addresses the concentration of the norm of linear transformations of sparse $\alpha$-subexponential random vectors. We anticipate that our new concentration inequalities will find further applications in the analysis of sketching algorithms [25] and for the large deviation principle for weighted random graphs [8].
@@
Organization: The rest of the paper is organized as follows. Section 2 reviews existing literature on sparse Hanson-Wright inequalities and concentration of sums of sparse random variables, along with a discussion and comparison of our findings with previous work. We present two applications of our sparse Hanson-Wright inequality in Section 3. Generalization of our main results to general matrix and non-centered random vectors are given in Section 4. The proofs of our main theoretical results are provided in Sections 5 and 6, while Section 7 contains the proofs of the applications.
Auxiliary results are presented in Appendices A. Appendices B and C provide additional proofs for the applications. Finally, the proof of corollaries are presented in Appendix D.
@@ 2 Previous results and comparison
Recent progress in the Hanson-Wright-type inequalities and more generally, the polynomials of independent random variables, can be found in [43, 60, 2, 67, 34, 65, 38, 11, 68, 46, 22, 32, 53, 8, 37, 55, 24, 53, 32, 19, 61] and references therein. For the following discussion in this section, let $A$ denote a diagonal-free, symmetric matrix. We mainly focus on the results of concentration for quadratic and linear forms of $\alpha$-subexponential random vectors and sparse random vectors.
Quadratic forms for $\alpha$-subexponential random vectors
Specifically, an extension from sub-gaussian to $\alpha$-subexponential random vectors was obtained in [32, Corollary 1.4 (2)]: for any $\alpha\in(0,1]$, let $X_{1},\dots,X_{n}$ be independent centered random variables with $\|X_{i}\|_{\Psi_{\alpha}}\leq K$. For any $t\geq 0$,
@@
The extension of (2.2) to any $\alpha\in(0,2]$ is discussed in Sambale’s work [53]. Note that for the results mentioned above, their sub-gaussian tail depends on the $\Psi_{\alpha}$-norm of the $X_{i}$, which does not capture the sharp dependence on the sparsity parameter $p$ when $X_{i}$ is sparse.
Quadratic forms for sparse random vectors
Now we turn to the results that have a better dependence on $p$ compared to (2.2) for sparse random vectors. The works [67, 51] have focused on quadratic forms involving sparse sub-gaussian random variables. Zhou [67] obtained a result analogous to (2.2) with $\alpha=2$, incorporating $p_{i}$’s only in the sub-gaussian tail, while [51] derived a comparable bound allowing for correlation. In a follow-up work, Zhou [68] examined quadratic forms of non-identical independent Bernoulli variables, bounding the moment generating function of the off-diagonal term. Augeri [8] studied sparse bounded random variables with uniform sparsity $p$ and $\|A\|\leq 1$, obtaining sharp tail bounds in the large deviation regime for $1\gg p\gg\log n/n$ (Lemma 2.8 from [8]). He, Knowles, and Marcozzi [34] established an upper bound on the $L_{r}$ norms $\|X^{T}AX\|_{L_{r}}$ for $X_{i}$’s centered $\mathrm{Ber}(p)$, which proved crucial for analyzing sparse Erdős-Rényi random graphs in the supercritical regime. Their result has been generalized to sparse sub-gaussian random vectors in [9, Lemma 7.4]. While our Theorems 4 and 5 draw inspiration from their work, they ultimately surpass the bounds established in [34].
The study of the concentration of U-statistics also offers insights that are useful for deriving tail estimates of random quadratic forms. In Giné, Latała, and Zine’s study [31], by applying decoupling results and performing straightforward calculations, one can obtain the following tail estimate as a special case of their Theorem 3.3 in [31]: consider $x=(x_{1},\cdots,x_{n})^{T}$ with independent components $x_{i}$ being centered Bernoulli random variables with parameter $p_{i}$. For any $t>0$,
where $B=(\max_{i}p_{i})\|A\|$ and $D=\max_{j}\sqrt{\sum_{i}a_{ij}^{2}p_{i}}$.
Recently, Chakrabortty and Kuchibhotla [20] studied a class of unbounded U-statistics. Their results (Theorem 1 in [20]) yield moment estimates of the quadratic forms for sparse $\alpha$-subexponential random variables with each $X_{i}=x_{i}\xi_{i}$, where $x_{i}\sim\mathrm{Ber}(p)$ and $\xi_{i}$ are independent, centered $\alpha$-subexponential random variables. Notably, for $\alpha\in(0,1]$, the tail bound implied by these moment estimates is consistent with (2) for constant $p$, up to the $\log n$ factors. For $\alpha>1$, the tail bound is equivalent to that of $\alpha=1$, again up to $\log n$ factors. Recent work by [10] has extended the results in [20], generalizing their findings to a broader class of unbounded, heavy-tailed U-statistics.
Schudy and Sviridenko [54] derived concentration inequalities for polynomials of independent central moment bounded random variables, improving upon their earlier work [55]. The central moment bounded random variables encompass bounded and continuous or discrete log-concave random variables. When $X_{1},\ldots,X_{n}$ are independent sparse exponential random variables, where $X_{i}=x_{i}\eta_{i}$ with independent $x_{i}\sim\mathrm{Ber}(p_{i})$ and centered $\eta_{i}\sim\mathrm{Exp}(K^{-1})$, their results yield:
@@
This result (2.4) can be generalized to sparse $\alpha$-subexponential random variables for all $\alpha\in[1,\infty]$, using a contraction principle (see Theorem 1 in [41]) alongside the decoupling technique.